# GENIAL: Generative Design Space Exploration via Network Inversion for Low Power Algorithmic Logic Units

### v1

@@ V-D1 Optimizing Finite State Machines
To evaluate the generalizability of our method, we applied it to Moore-type Finite State Machines (FSM), a common logic block beyond arithmetic units. We randomly generate a 16-state FSM (using 4-bit state encodings) with 3 inputs, 2 outputs, and 15 transition clauses. Starting from a random 30k-sample initial dataset and producing 5k encodings per GR with GENIAL, we observe that the minimum and mean NT are significantly lower than the initial dataset (GR0) (see Fig. 14).
V-D2 System-level
@@
While optimizing the operand encoding of a multiplier can reduce its power consumption, the impact must be assessed at the system level. Specifically, inputs to the multiplier, typically represented in TC, would need to be converted to a custom encoding used by the multiplier. This conversion is performed by an encoder (Enc.). Their size and cost are visible Table III (b). However, if the same operand can be converted once and reused multiple times, as is common in neural network accelerators, the cost of this conversion is amortized by the reuse factor. In fact, one could envision storing model weights directly in the custom encoding to further reduce system-wide latency and energy consumption.
VI Conclusion
@@
We introduced GENIAL, a machine-learning-driven framework for the automatic generation and optimization of combinational logic circuits. By combining a high-effort logic synthesis optimization flow with model inversion of a Transformer-based surrogate model for encoding generation, GENIAL enables sample-efficient design space exploration for optimizing key hardware metrics. Our results empirically demonstrate that two’s-complement and signed-magnitude encodings are both area- and power-efficient. Remarkably, GENIAL rediscovers these formats or close variants from scratch, without prior knowledge, and uncovers novel representations along the area-power Pareto frontier. Notably GR13, a 4b interger encoding resulting in 18% lower power consumption than the two’s complement. These findings validate surrogate-model inversion as a powerful strategy for circuit optimization. Future work will explore scaling GENIAL to larger circuits, other input value distributions, and circuits not as thoroughly investigated as multipliers, unlocking broader opportunities for logic function optimization.

### v2

@@ V-D1 Optimizing Finite State Machines
To evaluate the generalizability of our method, we applied it to Moore-type Finite State Machines (FSM), a common logic block beyond arithmetic units. We randomly generated a 16-state FSM (using 4-bit state encodings) with 3 inputs, 2 outputs, and 15 transition clauses. Starting from a random 30k-sample initial dataset and producing 5k encodings per GR with GENIAL, we observe that the minimum and mean NT are significantly lower than the initial dataset (GR0) (see Fig. 14).
Note that GENIAL is built for encoding discovery, and therefore cannot be applied to any random logic designs. GENIAL specifically targets problems with optimizable encodings.
V-D2 System-level considerations
@@
While optimizing the operand encoding of a multiplier can reduce its power consumption, the impact must be assessed at the system level. Specifically, inputs to the multiplier, typically represented in TC, would need to be converted to a custom encoding used by the multiplier. This conversion is performed by an encoder (Enc.). Their size and cost are visible in Table IV (b). However, if the same operand can be converted once and reused multiple times, as is common in neural network accelerators, the cost of this conversion is amortized by the reuse factor.
Finally, we validated the study by implementing a DaVinci-like core [cube], which realized a General Matrix Multiplication (GEMM) operation (cube) containing $16\times 32\times 16$ multipliers. We used commercial tools for synthesis, back-annotation for glitch-aware simulation, and power extraction. By swapping out the multipliers and placing the encoders right before the matrix buffers (L0), we observed major gains:
  1. i.
    SME multipliers enable up to 16% power savings against their two’s complement counterpart on a 2D convolution operation, where both operand matrices are filled with values randomly sampled from a normal distribution with $\sigma=3.0$.
  2. ii.
    By adding the encoders’ power to the total cost, this gain is reduced to 14%.
  3. iii.
    The total area of the GEMM operator increases by 5%.
In fact, one could envision storing model weights directly in the custom encoding to further reduce system-wide latency and energy consumption.
VI Limitations
While GENIAL demonstrates that encoding-based DSE can uncover novel arithmetic designs, several limitations remain:
VI-1 Runtime of high-fidelity flows
Our framework offers multiple flows with different fidelity/cost trade-offs. Complexity (CPX) and Yosys-based transistor count (NT) can be extracted very quickly, while high-effort synthesis optimization (FLOWY) provides more realistic estimates but requires significant runtime. This limits the size of the search space that can be explored at full fidelity. In practice, we balance speed and accuracy by mixing flows depending on the experiment, but scaling to larger multipliers will require further optimization (e.g., parallelization, lightweight proxies, or adaptive flow selection).
VI-2 LUT-based design representation
All candidate multipliers are generated as LUT circuits, rather than heuristics-based or manually designed RTL implementations. This abstraction was necessary to systematically explore arbitrary input encodings while ensuring that the output remained in two’s complement, enabling compatibility with existing systems. However, LUT-based RTL can yield suboptimal intermediate results (e.g., higher switching activity compared to hand-tuned RTL) before synthesis optimization. Future work could investigate direct RTL templates or hybrid representations to close this gap.
VI-3 Scale of benchmarks
Due to the runtime and design space explosion, our experiments focus on 4-bit multipliers and randomly generated FSMs. Although there are no standard benchmarks for encoding-driven arithmetic design, further validation on larger and more diverse circuits, such as 8- or 16-bit multipliers, square-root units, and other arithmetic operators, would better demonstrate the scalability and generality of our approach.
VI-4 Exploration capability of the design recommendation
While GENIAL discovers encodings that reduce switching activity, the current exploration did not yet reach or surpass the SME format. Enhancing the search capability of the design recommendation through improved diversity control or refined objectives may help uncover superior designs. Additionally, considering alternative constraints or scenarios (e.g., multi-modal data distributions) could reveal new encodings that outperform known formats.
VI-5 Single-objective exploration guidance
The current pipeline relies on a single proxy metric to provide an optimization signal through a surrogate model. Future extensions could incorporate multiple proxies, or even an uncertainty-aware exploration strategy, enabling active sampling in regions of high predictive uncertainty. This could reveal better local minima, improve search guidance, and enhance robustness across design spaces.
Together, these limitations highlight avenues for future work, including improving runtime efficiency, exploring richer and more direct RTL representations, enhancing the diversity and multi-objective guidance of the search, and scaling evaluations to larger, more complex arithmetic benchmarks.
VII Conclusion
@@
We introduced GENIAL, a machine-learning-driven framework for the automatic generation and optimization of combinational logic circuits. By combining a high-effort logic synthesis optimization flow with model inversion of a Transformer-based surrogate model for encoding generation, GENIAL enables sample-efficient design space exploration for optimizing key hardware metrics. Remarkably, GENIAL rediscovers the two’s-complement encoding and identifies close variants of the sign-magnitude format from scratch, without prior knowledge, while uncovering novel representations along the area–power Pareto frontier. Notably, GR13 is a 4-bit integer encoding resulting in 18% lower power consumption than the two’s complement. These findings validate surrogate-model inversion as a powerful strategy for circuit optimization. Future work will explore scaling GENIAL to larger circuits, other input value distributions, and circuits not as thoroughly investigated as multipliers, unlocking broader opportunities for logic function optimization.